Question 2 Total 8 Marks You are a management internee at the Gadgets Company. The company operates with two Divisions, Xen and Yen and produces designer toy cars. For Division Xen, the management applies overhead costs on the basis of direct labor cost, and for Division Yen, it applies overhead costs on the basis of machine-hours. The company completed Job FIN-E in late 2017. The companys books reveal the following additional information: Budgeted Amounts Labor cost (Direct) Manufacturing overhead Machine-hours Division Xen Division Yen $268 000 $285 000 51 400 mh $283 000 $328 000 61 300 mh Actual data for Tob FIN-E Division Xen Direct materials used Labor cost (Direct) Machine-hours $21 100 $21 500 9500 mhh Division Yen $26 300 $24 400 6300 mh Required: Your supervisor has asked you to work out the following, showing all your calculations (round off to two decimal places) a. The manufacturing overhead rate (budgeted) for Division Xen. b. The manufacturing overhead rate (budgeted) for Division Yen. c. The total overhead cost (budgeted) allocated to Job FIN-E. d. The unit cost for job FIN-E if the Job required Gadgets to produce 1065 designer tov cars.
Answer #1

a) Manufacturing overhead rate for division Xen = 328000*100/283000 = 115.90% of direct labour cost

b) Manufacturing overhead rate for division yen = 285000/51400 = 5.54 per machine hour

c) Total manufacturing cost allocated to job FIN-E = (21500*115.90%+6300*5.54) = 59820.50 or 59821

d) Unit cost = (21100+26300+21500+24400+59821)/1065 = 143.78 per unit